OpenStudy (jjuden):

Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 72 when x = 3.

OpenStudy (whpalmer4):

if \(y\) varies jointly with \(x^2\), then \(y = kx^2\). Substitute \(y = 72, x=3\) into that equation and solve for \(k\).

OpenStudy (anonymous):

@whpalmer4 so jointly and directly same thing?

OpenStudy (whpalmer4):

@timo86m jointly means you have something like \(z = k x y\) where \(z\) varies jointly with \(x\) and \(y\)

OpenStudy (jjuden):

y = 8x^2 y = 24x^2 y=1/8x^2 y=1/24x^2

OpenStudy (anonymous):

aaa but she said directly

OpenStudy (whpalmer4):

Yes, the answer is one of those :-)

OpenStudy (anonymous):

direct will be y=c*x^2 you are given y and x so solve for c :) like before 72=c*3^2

OpenStudy (whpalmer4):

\[y = k x^2\]\[y = 72, \, x = 3\]\[72 = k(3)^2\]Solve for \(k\)

OpenStudy (jjuden):

k=8 right?

OpenStudy (anonymous):

sorry os broke

OpenStudy (anonymous):

you are right

OpenStudy (whpalmer4):

@timo86m Sorry, yes, she did say directly, and I wrote jointly. The previous problem was a jointly, so I had jointly on my mind :-) Directly with \(x^2\) could be viewed as equivalent to jointly with \(x\) and \(x\), however! Aside from the incorrect word, everything else was correct.